### v2

# Enhanced Phase Estimation via Photon-Added Two-Mode Squeezed States and Kerr Nonlinearity

Abstract
Quantum metrology employs quantum resources to achieve measurement precision beyond classical limits. This work investigates a Mach–Zehnder interferometer incorporating a Kerr nonlinear phase shifter, with photon-added two-mode squeezed coherent states generated via four-wave mixing as input. We demonstrate that increasing both the photon-addition order and the input resource strength systematically enhances phase sensitivity, quantum Fisher information, and the corresponding quantum Cramér–Rao bound. The proposed system not only surpasses the standard quantum limit but also approaches or exceeds the Heisenberg limit for linear phase shifts, while Kerr nonlinearity enables surpassing the super-Heisenberg limit. Furthermore, the scheme exhibits enhanced robustness against photon loss, providing a promising pathway toward practical high-precision quantum metrology applications.
PACS: 03.67.-a, 05.30.-d, 42.50.Dv, 03.65.Wj
I Introduction
Precision measurement constitutes the cornerstone of scientific and technological advancement, with its fundamental limits dictating our ability to probe and manipulate physical phenomena. From traditional mechanical manufacturing to modern lithography, and from gravitational-wave detection to biomedical sensing, progress consistently relies on advancing measurement precision. Within classical frameworks, measurement accuracy is fundamentally constrained by the standard quantum limit (SQL), where measurement uncertainty scales as $1/\sqrt{\bar{N}}$ with the average photon number $\bar{N}$ inside the interferometer 1 ; 2 ; 3 .
Quantum mechanics offers transformative strategies to overcome these limitations. Quantum metrology exploits distinct quantum resources—including superposition, entanglement, and nonclassical correlations—to surpass the SQL and achieve ultra-high precision approaching or exceeding the Heisenberg limit (HL), which scales as $1/\bar{N}$ 4 ; 5 ; 6 ; 7 ; 8 ; 9 ; 10 ; 11 . This field carries profound physical implications and substantial potential for applications ranging from gravitational-wave detection and atomic clocks to magnetometry, super-resolution imaging, and measurements of fundamental physical constants.
A primary strategy for quantum enhancement involves injecting nonclassical states into interferometric devices rather than conventional coherent states. Squeezed states and entangled states generated via parametric amplification processes have proven effective in surpassing the SQL. However, preparing ideal entangled states (e.g., NOON states) remains challenging due to their extreme susceptibility to environmental noise and photon loss, which significantly limits practical implementation 11 ; 12 ; 13 ; 14 ; 15 ; 16 .
Recent research has extensively explored non-Gaussian operations, particularly photon addition and subtraction, which can substantially enhance the nonclassical properties of quantum states and potentially improve measurement robustness. Among these, photon-addition operations offer distinct advantages in quantum-state preparation and in enhancing both nonclassical characteristics and measurement precision 17 ; 18 ; 19 ; 20 ; 21 ; 22 ; 23 ; 24 ; 25 ; 26 ; 27 ; 28 ; 29 ; 30 ; 31 ; 32 . Concurrently, interferometer architectures have been continuously refined. For instance, incorporating Kerr nonlinear media to replace conventional linear phase shifters can theoretically achieve accuracies beyond the sub-HL ($1/\bar{N}^{3/2}$), potentially reaching the super-Heisenberg limit (SHL), i.e., $1/\bar{N}^{2}$ 33 ; 34 ; 35 ; 36 ; 37 ; 38 ; 39 ; 40 ; 41 .
Building upon this foundation, this work develops a quantum precision measurement scheme that combines high precision with enhanced robustness. We generate two-mode squeezed coherent states (TMSCS) via four-wave mixing (FWM) and apply photon-addition operations to create photon-added TMSCS (PA-TMSCS) as the interferometer input resource. Furthermore, we replace the linear phase shifter in a conventional Mach–Zehnder interferometer (MZI) with a Kerr nonlinear phase shifter, constructing a Kerr nonlinear MZI (KMZI). This study systematically analyzes the scheme’s performance under both ideal conditions and in the presence of photon loss. By examining phase sensitivity via intensity-difference detection and quantum Fisher information (QFI), and comparing linear and nonlinear phase-shift scenarios, we verify the synergistic advantages of photon-addition operations and Kerr nonlinear effects in enhancing measurement accuracy and system resilience. This work provides valuable theoretical insights for developing practical quantum-enhanced measurement technologies.
The paper is organized as follows: Section II introduces the phase estimation model based on a KMZI with PA-TMSCS input. Section III investigates the phase sensitivity of intensity-difference detection for both linear ($k=1$) and Kerr nonlinear ($k=2$) phase shifts, including the effects of photon loss. Section IV examines the QFI for these cases under ideal and lossy conditions. Conclusions are presented in the final section.
II Phase Estimation Model Based on a KMZI with PA-TMSCS Input
Figure 1: Schematic of phase estimation using PA-TMSCS inputs in a KMZI. Photon-addition operations are applied to both the $a$-mode and $b$-mode of the TMSCS. Photon loss inside the KMZI is simulated by a virtual beam splitter, and intensity-difference detection is implemented at the output port.
This section presents the phase estimation model employing PA-TMSCS inputs in a KMZI. As illustrated in Fig. 1, TMSCS can be generated by applying an FWM process induced by a pump field to coherent states. The FWM process is equivalent to a two-mode squeezing operation described by the operator $S_{2}(r)=\exp[\xi^{\ast}ab-\xi a^{\dagger}b^{\dagger}]$, where $\xi=re^{i\phi_{1}}$, with $r$ and $\phi_{1}$ representing the squeezing parameter and phase, respectively. To simplify experimental procedures and avoid complicated phase-matching conditions, only the $a$-mode input is prepared as a coherent state $\lvert\alpha\rangle_{a}$, while the $b$-mode input remains in the vacuum state $\lvert 0\rangle_{b}$. The coherent state satisfies $a\lvert\alpha\rangle_{a}=\alpha\lvert\alpha\rangle_{a}$, where $\alpha=\lvert\alpha\rvert e^{i\theta}$ ($\lvert\alpha\rvert$ and $\theta$ denote amplitude and phase, respectively). For simplicity and to satisfy phase-matching conditions, we set $\theta=0$ ($\alpha=\lvert\alpha\rvert$) and $\phi_{1}=\pi$, which proves advantageous for enhanced phase estimation 24 ; 42 .
To further optimize the input states for improved measurement accuracy, we implement non-Gaussian photon-addition operations on the TMSCS. The resulting PA-TMSCS serves as the input state and can be expressed as
$\lvert\text{in}\rangle=\frac{1}{\sqrt{N_{m,n}}}a^{\dagger m}b^{\dagger n}S_{2}(r)\lvert\alpha\rangle_{a}\lvert 0\rangle_{b},$ (1)
where $a^{\dagger m}$ and $b^{\dagger n}$ represent photon-addition operations on the $a$-mode and $b$-mode, respectively, and $N_{m,n}$ is the normalization coefficient. To facilitate calculations, the photon-addition operators in Eq. (1) can be converted into partial-derivative form:
$\lvert\text{in}\rangle$ $=$ $\frac{1}{\sqrt{N_{m,n}}}\frac{\partial^{m+n}}{\partial t_{1}^{m}\partial\tau_{1}^{n}}\exp[a^{\dagger}t_{1}]$ (2)
$\times\exp[b^{\dagger}\tau_{1}]S_{2}(r)\lvert\alpha\rangle_{a}\lvert 0\rangle_{b}\Big|_{t_{1}=\tau_{1}=0}.$
Based on this representation and quantum-state normalization, the expression for $N_{m,n}$ can be derived through relevant operator transformation relations:
$N_{m,n}=\Gamma_{0}\Bigl\{$ $\exp\bigl[((t_{1}t_{2}+\tau_{1}\tau_{2})\cosh r$ (3)
$+(t_{1}\tau_{1}+t_{2}\tau_{2})\sinh r)\cosh r\bigr]$
$\times\exp\bigl[((t_{1}+t_{2})\cosh r$
$+(\tau_{1}+\tau_{2})\sinh r)\alpha\bigr]\Bigr\},$
where
$\Gamma_{0}\{\cdot\}=\frac{\partial^{2m+2n}}{\partial t_{1}^{m}\partial\tau_{1}^{n}\partial t_{2}^{m}\partial\tau_{2}^{n}}\{\cdot\}\Big|_{t_{1}=\tau_{1}=t_{2}=\tau_{2}=0}.$ (4)
Conventional MZIs constitute linear measuring instruments comprising two 50:50 beam splitters (BS1 and BS2) and a linear phase shifter. Following Yurke et al. 43 , using angular momentum operators in the Schwinger representation, the equivalent operators for BS1 and BS2 are $B_{1}=\exp[-i\pi(a^{\dagger}b+ab^{\dagger})/4]$ and $B_{2}=\exp[i\pi(a^{\dagger}b+ab^{\dagger})/4]$, respectively. These operators satisfy the transformation relations:
$B_{1}^{\dagger}\binom{a}{b}B_{1}=\frac{\sqrt{2}}{2}\begin{pmatrix}1&-i\\ -i&1\end{pmatrix}\binom{a}{b}$, (5)
and
$B_{2}^{\dagger}\binom{a}{b}B_{2}=\frac{\sqrt{2}}{2}\begin{pmatrix}1&i\\ i&1\end{pmatrix}\binom{a}{b}$. (6)
For our investigation, we replace the linear phase shifter in the conventional MZI with a Kerr nonlinear phase shifter to construct an improved KMZI. The phase shifter’s equivalent operator is $U(\phi,k)=\exp[i\phi(a^{\dagger}a)^{k}]$, where $k=1$ corresponds to a linear phase shift and $k=2$ to a Kerr nonlinear phase shift. Here, $\phi$ represents the phase shift to be measured. The transformation relations for linear and Kerr nonlinear phase shifts are respectively given by:
$U^{\dagger}(\phi,1)aU(\phi,1)=e^{i\phi}a,$ (7)
and
$U^{\dagger}(\phi,2)aU(\phi,2)=e^{i\phi}e^{i\phi(2a^{\dagger}a)}a.$ (8)
The transformation relation for $U(\phi,2)$ in Eq. (8) can be derived using methods from Ref. 40 based on Fock-state representation, with detailed steps provided in Appendix A.
During actual measurements, photon losses inevitably occur. In our model (Fig. 1), we simulate photon losses in the KMZI’s internal $a$-mode using a virtual beam splitter placed between BS1 and the phase shifter. The corresponding transformation is:
$B_{f}^{\dagger}\begin{pmatrix}a\\ a_{f}\end{pmatrix}B_{f}=\begin{pmatrix}\sqrt{1-l}&\sqrt{l}\\ -\sqrt{l}&\sqrt{1-l}\end{pmatrix}\begin{pmatrix}a\\ a_{f}\end{pmatrix}$, (9)
where $a_{f}$ is the photon annihilation operator for the auxiliary mode $f$ containing vacuum noise $\lvert 0\rangle_{f}$, while $B_{f}$ and $l$ represent the equivalent operator and reflectivity of the virtual beam splitter, respectively. The parameter $l$ corresponds to the photon loss rate, satisfying $0\leq l\leq 1$. Higher $l$ values indicate increased loss, with $l=0$ representing the lossless case and $l=1$ corresponding to complete absorption.
Within this KMZI framework, intensity-difference detection is implemented by placing photon counters at the output ports. Since intensity-difference detection is relatively straightforward to implement experimentally, it finds widespread use for phase sensitivity measurements in MZIs.
III Phase Sensitivity of Intensity-Difference Detection
The phase sensitivity for intensity-difference detection can be expressed using the error propagation formula:
$\Delta\phi=\frac{\sqrt{\langle I_{D}^{2}\rangle-\langle I_{D}\rangle^{2}}}{\bigl\lvert\frac{\partial\langle I_{D}\rangle}{\partial\phi}\bigr\rvert}$, (10)
where $I_{D}=a^{\dagger}a-b^{\dagger}b$, and $\langle\cdot\rangle=\langle\text{out}\rvert\cdot\lvert\text{out}\rangle$ denotes the expectation value with respect to the output state. From the phase estimation model description, the output state can be expressed as:
$\lvert\text{out}\rangle=B_{2}U(\phi,k)B_{f}B_{1}\lvert\text{in}\rangle\lvert 0\rangle_{f}.$ (11)
We focus on phase sensitivity for both $k=1$ and $k=2$ cases. For brevity, explicit expressions for phase sensitivity are omitted here. Substituting the expectation values $\langle I_{D}\rangle$ and $\langle I_{D}^{2}\rangle$ into Eq. (10) yields the phase sensitivity. The specific expressions for $\langle I_{D}\rangle$ and $\langle I_{D}^{2}\rangle$ under $k=1$ and $k=2$ conditions are provided in Appendices B and C, respectively.
Figure 2: Phase sensitivity $\Delta\phi$ versus phase shift $\phi$ for squeezing parameter $r=0.9$ and coherent amplitude $\alpha=1$. (a) and (b): linear phase shift ($k=1$); (c) and (d): Kerr nonlinear phase shift ($k=2$). The black solid line corresponds to $m=n=0$ (no photon addition).
Fig. 2 illustrates the variation of phase sensitivity $\Delta\phi$ with phase shift $\phi$ for $k=1$ and $k=2$ cases, where Figs. 2(a) and (b) depict linear phase shift ($k=1$), while Figs. 2(c) and (d) represent Kerr nonlinear phase shift ($k=2$). According to our computational expressions and the behavior of $\Delta\phi$ in Fig. 2(a), the variation period is $\pi$. Comparing Figs. 2(a) and (c), curves incorporating photon-addition operations show significant improvement over the black solid line ($m=n=0$). Thus, for schemes employing PA-TMSCS input states, photon addition effectively enhances phase sensitivity. Notably, the phase sensitivity variation range in Fig. 2(c) is substantially reduced compared to Fig. 2(a) due to the Kerr nonlinear effect at $k=2$ (hence Fig. 2(c) is not plotted over a full period), suggesting that Kerr nonlinear phase shifting can significantly improve phase sensitivity within certain phase shift ranges.
To further investigate effective photon-addition schemes, we analyze phase sensitivity characteristics under narrowed plotting ranges, as shown in Figs. 2(b) and (d). Phase sensitivity improves with increasing photon-addition order $m+n$, and near the optimal phase sensitivity point, symmetric photon addition ($m=n$) yields the best performance for a given $m+n$. This indicates that performing symmetric photon-addition operations in both $a$- and $b$-modes is more conducive to enhancing measurement accuracy. Consequently, we focus on the $m=n$ scenario for photon-addition operations throughout this paper.
Figure 3: For fixed phase shift $\phi=3.12$: (a) phase sensitivity $\Delta\phi$ versus squeezing parameter $r$ with coherent amplitude $\alpha=1$; (b) $\Delta\phi$ versus $\alpha$ with $r=0.9$. Solid (dashed) lines correspond to $k=2$ ($k=1$).
Fig. 3 demonstrates that phase sensitivity $\Delta\phi$ improves with increasing squeezing parameter $r$, coherent amplitude $\alpha$, and photon-addition order. Additionally, comparing solid and dashed lines clearly shows that Kerr nonlinear phase shift provides improved phase sensitivity compared to linear phase shift.
In practical scenarios, photon loss constitutes an inevitable factor. We subsequently analyze its impact on phase sensitivity and compare achievable measurement precision limits. The measurement precision limit can be defined through the average photon number inside the interferometer $\bar{N}$, expressed as:
$\bar{N}$ $=$ ${}_{\text{int}}\langle\psi\rvert a^{\dagger}a+b^{\dagger}b\lvert\psi\rangle_{\text{int}}$ (12)
$=$ $D_{1}(1,0,1,0)+D_{1}(0,1,0,1),$
where $\lvert\psi\rangle_{\text{int}}=U(\phi,k)B_{f}B_{1}\lvert\text{in}\rangle\lvert 0\rangle_{f}$, and specific expressions for $D_{1}(1,0,1,0)$ and $D_{1}(0,1,0,1)$ are given in Eq. (B5) of Appendix B.
Figure 4: Linear phase shift case ($k=1$): phase sensitivity $\Delta\phi$ versus loss rate $l$ for (a) $m=n=0$, (b) $m=n=1$, (c) $m=n=2$, with fixed phase shift $\phi=3.12$, squeezing parameter $r=0.9$, and coherent amplitude $\alpha=1$. SQL and HL are shown for comparison.
Figure 5: Kerr nonlinear phase shift case ($k=2$): phase sensitivity $\Delta\phi$ versus loss rate $l$ for (a) $m=n=0$, (b) $m=n=1$, (c) $m=n=2$, with fixed phase shift $\phi=3.12$, squeezing parameter $r=0.9$, and coherent amplitude $\alpha=1$. SQL, HL, sub-HL, and SHL are shown for comparison.
Figs. 4(a)–(c) depict linear phase shift sensitivity $\Delta\phi$ as a function of loss rate $l$, compared with SQL and HL. Despite photon loss, phase sensitivity can still surpass SQL, with photon addition improving system robustness. Figs. 5(a)–(c) further illustrate Kerr nonlinear phase shift sensitivity versus loss rate $l$. Comparison with Fig. 4 clearly shows that Kerr nonlinearity enhances phase sensitivity by an order of magnitude, enabling exceedance of the HL threshold even with loss, whereas linear phase shift only surpasses SQL within limited $l$ ranges. However, with fixed $r=0.9$ and $\alpha=1$, phase sensitivity for $m=n=0$ can exceed sub-HL, but not for $m=n=1,2$. This may arise from Kerr nonlinearity sensitivity to other parameters (e.g., phase shift). This situation indicates that without photon addition, phase sensitivity can exceed sub-HL within certain ranges, which is relatively favorable for Kerr nonlinearity.
IV Quantum Fisher Information
IV.1 QFI of KMZI in Ideal Case
Quantum Fisher information (QFI) quantifies the maximum information about phase shift $\phi$ that a quantum state can convey after passing through a phase shifter. In our scheme, employing linear and Kerr nonlinear phase shifters within the MZI leads to two corresponding QFI representations, denoted $F_{k}$ ($k=1,2$). QFI is independent of any specific detection scheme and provides an upper bound on classical Fisher information obtainable through $\phi$ estimation. The optimal error bound for phase sensitivity, known as the quantum Cramér–Rao bound (QCRB), is given by 44 ; 45 ; 46 :
$\Delta\phi_{\text{QCRB}_{k}}=\frac{1}{\sqrt{F_{k}}}.$ (13)
For pure states, QFI under ideal conditions ($l=0$) can be calculated as 47 ; 48 :
$F_{k}=4\bigl[\langle\psi_{\phi}^{\prime}|\psi_{\phi}^{\prime}\rangle-\lvert\langle\psi_{\phi}^{\prime}|\psi_{\phi}\rangle\rvert^{2}\bigr],$ (14)
where $\lvert\psi_{\phi}\rangle=U(\phi,k)B_{1}\lvert\psi\rangle_{\text{in}}$ represents the quantum state before BS2 in the lossless KMZI, and $\lvert\psi_{\phi}^{\prime}\rangle=\partial\lvert\psi_{\phi}\rangle/\partial\phi$. For both linear ($k=1$) and Kerr nonlinear ($k=2$) phase shifts, QFI can be further simplified as 40 :
$F_{k}=4\langle\Delta^{2}n_{a}^{k}\rangle=4\bigl[\langle(n_{a}^{k})^{2}\rangle-\langle n_{a}^{k}\rangle^{2}\bigr],$ (15)
where $\langle\cdot\rangle=\langle\Psi\rvert\cdot\lvert\Psi\rangle$, with $\lvert\Psi\rangle=B_{1}\lvert\text{in}\rangle$. Here, $n_{a}^{k}=(a^{\dagger}a)^{k}$ ($k=1,2$). To facilitate QFI computation, the normal ordering form of $n_{a}^{w}$ is obtained via the operator identity:
$n_{a}^{w}$ $=$ $\frac{\partial^{w}}{\partial x^{w}}\exp\bigl[xa^{\dagger}a\bigr]\Big|_{x=0}$ (16)
$=$ $\frac{\partial^{w}}{\partial x^{w}}\colon\!\exp\bigl[(e^{x}-1)a^{\dagger}a\bigr]\!\colon\Big|_{x=0}.$
Since $a^{\dagger}a$ commutes with the phase shifter operator and losses are neglected here, utilizing the normal ordering of $n_{a}^{w}$ allows deriving the QFI for $k=1$ using Eq. (B5) with $l=0$:
$F_{1}=4\bigl[D_{1}(2,0,2,0)+D_{1}(1,0,1,0)-D_{1}(1,0,1,0)^{2}\bigr],$ (17)
while QFI for $k=2$ becomes:
$F_{2}$ $=$ $4\bigl[D_{1}(4,0,4,0)+6D_{1}(3,0,3,0)$ (18)
$+7D_{1}(2,0,2,0)+D_{1}(1,0,1,0)$
$-\bigl(D_{1}(2,0,2,0)+D_{1}(1,0,1,0)\bigr)^{2}\bigr].$
Figure 6: For $k=1$ case: (a) QFI $F_{1}$ versus squeezing parameter $r$ with coherent amplitude $\alpha=1$; (b) $F_{1}$ versus $\alpha$ with $r=0.9$. For $k=2$ case: (c) QFI $F_{2}$ versus $r$ with $\alpha=1$; (d) $F_{2}$ versus $\alpha$ with $r=0.9$.
As evident from Fig. 6, increasing photon-addition order significantly enhances QFI. Furthermore, comparing Figs. 6(a)–(b) with Figs. 6(c)–(d) clearly shows that the $k=2$ case yields larger QFI values. This demonstrates that both photon-addition operations and Kerr nonlinear phase shifts effectively improve QFI.
IV.2 Impacts of Photon Losses on QFI
Figure 7: Theoretical model of lossy KMZI using PA-TMSCS as input. Two virtual beam splitters, positioned before and after the phase shifter, simulate photon losses. Parameter $l$ represents beam splitter reflectivity (loss rate), while $b_{v}$ denotes the vacuum operator.
IV.2.1 Effects of Photon Loss on QFI in Linear Phase Shift Case ($k=1$)
Practical measurements involve photon losses due to vacuum noise, affecting QFI. For the initial pure state $\lvert\Psi_{S}\rangle$ ($\lvert\Psi_{S}\rangle=B_{1}\lvert\psi\rangle_{\text{in}}$) in a probe system $S$ with a lossy MZI, we introduce environment $E$ orthogonal states $\lvert j_{E}\rangle$ and Kraus operators $\hat{\Pi}_{j}(\phi)$ to characterize $\lvert\Psi_{S}\rangle$ evolution through the phase shifter with photon losses. Fig. 7 illustrates the photon loss within the MZI, providing a visual representation of the model structure under investigation. Following Escher et al.’s methods for calculating QFI in open quantum systems 49 , the quantum states $\lvert\Psi_{S}\rangle$ and environment vacuum noise $\lvert 0_{E}\rangle$, after unitary evolution $U_{S+E}(\phi)$, can be expressed in the extended $S+E$ space:
$\lvert\Psi_{S+E}\rangle$ $=$ $U_{S+E}(\phi)\lvert\Psi_{S}\rangle\lvert 0_{E}\rangle$ (19)
$=$ $\sum_{j=0}^{\infty}\hat{\Pi}_{j}(\phi)\lvert\Psi_{S}\rangle\lvert j_{E}\rangle.$
Although photon losses transform $\lvert\Psi_{S}\rangle$ into a mixed state, this approach allows treating $\lvert\Psi_{S+E}\rangle$ in $S+E$ as a pure state. For the entire purified system, QFI under photon losses can be expressed as:
$F_{L_{1}}\leqslant C_{Q}\bigl[\lvert\Psi_{S}\rangle,\hat{\Pi}_{j}(\phi)\bigr]=4\bigl[\langle\hat{H}_{1}\rangle_{S}-\lvert\langle\hat{H}_{2}\rangle_{S}\rvert^{2}\bigr],$ (20)
where the lower bound of $C_{Q}[\lvert\Psi_{S}\rangle,\hat{\Pi}_{j}(\phi)]$ is shown to be the QFI for the reduced system 49 , and $\hat{H}_{1,2}$ are Hermitian operators defined as:
$\hat{H}_{1}$ $=$ $\sum_{j=0}^{\infty}\frac{d\hat{\Pi}_{j}^{\dagger}(\phi)}{d\phi}\frac{d\hat{\Pi}_{j}(\phi)}{d\phi},$ (21)
$\hat{H}_{2}$ $=$ $i\sum_{j=0}^{\infty}\frac{d\hat{\Pi}_{j}^{\dagger}(\phi)}{d\phi}\hat{\Pi}_{j}(\phi),$ (22)
with $\langle\cdot\rangle_{S}$ denoting expectation value $\langle\Psi_{S}\rvert\cdot\lvert\Psi_{S}\rangle$, representing the average of $\hat{H}_{1,2}$. Thus, $F_{L_{1}}=C_{Q\min}$, and for linear phase shift ($k=1$), the Kraus operator is:
$\hat{\Pi}_{j}(\phi)=\sqrt{\frac{l^{j}}{j!}}e^{i\phi(a^{\dagger}a-\gamma j)}(1-l)^{\frac{a^{\dagger}a}{2}}a^{j},$ (23)
where $l$ is the loss rate, and $\gamma=0$ or 1-1 corresponds to photon losses before or after the linear phase shifter, respectively (Fig. 7). Optimizing $\gamma$ yields $C_{Q\min}$. Thus, QFI with photon losses becomes 49 :
$F_{L_{1}}=\frac{4F_{1}(1-l)\langle n_{a}\rangle}{lF_{1}+4(1-l)\langle n_{a}\rangle}.$ (24)
IV.2.2 Effects of Photon Loss on QFI in Kerr Nonlinear Phase Shift Case ($k=2$)
Extending the previous approach, we investigate QFI under photon losses for Kerr nonlinear phase shift ($k=2$). The general Kraus operator form, incorporating Kerr nonlinear phase shift, is defined as 40 :
$\hat{\Pi}_{j}(\phi)=\sqrt{\frac{l^{j}}{j!}}e^{i\phi[(a^{\dagger}a)^{2}-2\mu_{1}a^{\dagger}aj-\mu_{2}j^{2}]}(1-l)^{\frac{a^{\dagger}a}{2}}a^{j},$ (25)
where parameters $\mu_{1}=\mu_{2}=0$ or 1-1 correspond to photon losses before or after the Kerr nonlinear phase shifter (Fig. 7). Referring to Eq. (20), we derive (see Appendix D for derivation):
$F_{L_{2}}$ $\leqslant$ $C_{Q}\bigl[\lvert\Psi_{S}\rangle,\hat{\Pi}_{j}(\phi)\bigr]$ (26)
$=$ $4\bigl[K_{1}^{2}\langle\Delta^{2}n_{a}^{2}\rangle_{S}-K_{2}\langle n_{a}^{3}\rangle_{S}$
$+K_{3}\langle n_{a}^{2}\rangle_{S}-K_{4}\langle n_{a}\rangle_{S}$
$-K_{5}\langle n_{a}^{2}\rangle_{S}\langle n_{a}\rangle_{S}-K_{6}\langle n_{a}\rangle_{S}^{2}\bigr],$
where $K_{i}$ ($i=1,2,3,4,5,6$) are detailed in Appendix D. To determine QFI $F_{QL}$ under photon losses for Kerr nonlinear phase shift ($k=2$), substitute optimal values $\mu_{1\text{opt}}$ and $\mu_{2\text{opt}}$ into $C_{Q}$ to find $C_{Q\min}$. Specific expressions for $\mu_{1\text{opt}}$ and $\mu_{2\text{opt}}$ are provided in Appendix D. Similar to Eq. (13), $\Delta\phi_{\text{QBL}_{k}}$ under photon losses for $k=1$ and $k=2$ is derived from Eqs. (24) and (26), respectively, as $\Delta\phi_{\text{QBL}_{k}}=1/\sqrt{F_{L_{k}}}$.
Figure 8: For $k=1$ case: QCRB $\Delta\phi_{\text{QBL}_{1}}$ versus loss rate $l$ for (a) $m=n=0$, (b) $m=n=1$, (c) $m=n=2$, with fixed squeezing parameter $r=0.9$ and coherent amplitude $\alpha=1$. SQL and HL are shown for comparison.
As shown in Fig. 8, despite photon loss, QCRB can still be improved through photon addition and can surpass SQL across wide loss rate $l$ ranges, even reaching HL for small $l$.
Figure 9: For $k=2$ case: QCRB $\Delta\phi_{\text{QBL}_{2}}$ versus loss rate $l$ for (a) $m=n=0$, (b) $m=n=1$, (c) $m=n=2$, with fixed squeezing parameter $r=0.9$ and coherent amplitude $\alpha=1$. SQL, HL, sub-HL, and SHL are shown for comparison.
Comparing Fig. 8 with Fig. 9 reveals that Kerr nonlinear phase shift ($k=2$) significantly enhances QCRB. Despite losses, QCRB improvement through photon addition persists, with QCRB breaking through HL and sub-HL across wide $l$ ranges, even exceeding SHL for small $l$. This demonstrates that although photon loss degrades QCRB, both photon addition and Kerr nonlinear phase shift can mitigate this degradation, thereby enhancing system robustness.
V Conclusion
We have proposed and analyzed an enhanced phase estimation scheme combining PA-TMSCS with a KMZI. Our results demonstrate that this approach systematically improves phase sensitivity and QFI, enabling the surpassing of fundamental precision limits. Specifically, the scheme exceeds the SQL and approaches the HL for linear phase shifts, while the Kerr nonlinearity facilitates scaling beyond the SHL. Moreover, the integrated design exhibits enhanced robustness against photon loss, a critical attribute for practical implementations.
This work elucidates the synergistic mechanism between non-Gaussian photon addition and Kerr nonlinearity, which concurrently boosts measurement precision and resilience. The findings establish a valuable theoretical framework for developing practical quantum-enhanced metrology protocols, with potential applications in ultra-precise sensing. Future work may focus on identifying optimal operating parameters and experimental realizations in physical platforms suitable for nonlinear quantum optics.
Acknowledgments
This work is supported by the National Natural Science Foundation of China (Grants No. 12564049 and No. 12104195) and the Jiangxi Provincial Natural Science Foundation (Grants No. 20242BAB26009 and 20232BAB211033), Jiangxi Provincial Key Laboratory of Advanced Electronic Materials and Devices (Grant No. 2024SSY03011), Jiangxi Civil-Military Integration Research Institute (Grant No. 2024JXRH0Y07), as well as the Science and Technology Project of Jiangxi Provincial Department of Education (Grant No. GJJ2404102).
Appendix A Derivation of Transformation Relation for Kerr Nonlinear Phase Shifter Operator
In the scheme proposed in this paper, the Kerr nonlinear phase shifter significantly enhances phase measurement accuracy. The associated calculations can be efficiently performed using the transformation relation of the equivalent operator of the Kerr nonlinear phase shifter, $U(\phi,2)=e^{i\phi(a^{\dagger}a)^{2}}$, as shown in our study. Therefore, this appendix provides a detailed derivation of the transformation relation $U^{\dagger}(\phi,2)aU(\phi,2)=e^{i\phi}e^{i\phi(2a^{\dagger}a)}a$ in Eq. (8). Here, we assume the operator corresponding to the transformation relation of $U(\phi,2)$ is $O$, i.e., $O=U^{\dagger}(\phi,2)aU(\phi,2)$. The matrix of $O$ in the Fock state space can be represented by $O_{n,n^{\prime}}=\langle n\rvert O\lvert n^{\prime}\rangle$, and its corresponding expansion is given by
$O=\sum_{n,n^{\prime}=0}^{\infty}\lvert n\rangle\langle n^{\prime}\rvert O_{n,n^{\prime}}.$ (A1)
By utilizing $\langle n\rvert a^{\dagger}a=\langle n\rvert n$ and $\langle n\rvert a=\langle n+1\rvert\sqrt{n+1}$, the matrix element $O_{n,n^{\prime}}$ in the aforementioned equation can be further calculated as
$O_{n,n^{\prime}}=\sqrt{n+1}e^{i\phi(2n+1)}\delta_{n+1,n^{\prime}},$ (A2)
and by substituting it into Eq. (A1), the transformation relation of $U(\phi,2)$ can be deduced as follows
$O$ $=e^{i\phi}\sum_{n=0}^{\infty}\lvert n\rangle\langle n+1\rvert\sqrt{n+1}e^{i2n\phi}$
$=e^{i\phi}\sum_{n=0}^{\infty}\lvert n\rangle\langle n\rvert e^{i2n\phi}a=e^{i\phi}e^{i\phi(2a^{\dagger}a)}a,$ (A3)
where we have used the completeness of the Fock state $\sum_{n=0}^{\infty}\lvert n\rangle\langle n\rvert=1$.
Appendix B Specific Derivations and Expressions for $\langle I_{D}\rangle$ and $\langle I_{D}^{2}\rangle$ under $k=1$ Condition
In order to obtain the mean values $\langle I_{D}\rangle$ and $\langle I_{D}^{2}\rangle$ in Eq. (10) in a simple way, the general formula is introduced, i.e.,
${}_{f}\langle 0\rvert\langle\text{in}\rvert B_{1}^{\dagger}B_{f}^{\dagger}e^{-i\phi a^{\dagger}a}$
$\times\bigl(a^{\dagger m_{1}}b^{\dagger n_{1}}a^{m_{2}}b^{n_{2}}\bigr)e^{i\phi a^{\dagger}a}B_{f}B_{1}\lvert\text{in}\rangle\lvert 0\rangle_{f}$
$=D_{1}(m_{1},n_{1},m_{2},n_{2}).$ (B1)
The quantum state used in the above formula to calculate the average value is the state after passing through the phase shifter, and the output state still needs to pass through BS2. Thus using the transformation relation Eqs. (6) and (B1) one can obtain $\langle I_{D}\rangle$ and $\langle I_{D}^{2}\rangle$ respectively as
$\langle I_{D}\rangle=i\bigl[D_{1}(1,0,0,1)-D_{1}(0,1,1,0)\bigr],$ (B2)
and
$\langle I_{D}^{2}\rangle$ $=D_{1}(0,1,0,1)+D_{1}(1,0,1,0)+2D_{1}(1,1,1,1)$
$\quad-D_{1}(2,0,0,2)-D_{1}(0,2,2,0).$ (B3)
In order to facilitate the derivation of general formula Eq. (B1), we use the following partial derivative form for representation,
$D_{1}(m_{1},n_{1},m_{2},n_{2})$
$={}_{f}\langle 0\rvert\langle\text{in}\rvert B_{1}^{\dagger}B_{f}^{\dagger}e^{-i\phi a^{\dagger}a}\frac{\partial^{m_{1}+n_{1}+m_{2}+n_{2}}}{\partial x_{1}^{m_{1}}\partial y_{1}^{n_{1}}\partial x_{2}^{m_{2}}\partial y_{2}^{n_{2}}}$
$\quad\exp\bigl[a^{\dagger}x_{1}+b^{\dagger}y_{1}+ax_{2}+by_{2}\bigr]\Big|_{x_{1}=y_{1}=x_{2}=y_{2}=0}$
$\quad\times e^{i\phi a^{\dagger}a}B_{f}B_{1}\lvert\text{in}\rangle\lvert 0\rangle_{f}.$ (B4)
According to the transformation relationship of the corresponding operator, it can be calculated that
$\Gamma_{1}\Bigl\{\exp\bigl[\alpha(t_{1}+t_{2}+A_{1}+A_{4})\cosh r\bigr]$
$\times\exp\bigl[\alpha(\tau_{1}+\tau_{2}+A_{2}+A_{3})\sinh r\bigr]$
$\times\exp\bigl[(t_{1}\sinh r+A_{2}\cosh r)\tau_{1}\cosh r$
$\quad+(\tau_{2}\sinh r+A_{1}\cosh r)t_{2}\cosh r\bigr]$
$\times\exp\bigl[(t_{1}\cosh r+A_{2}\sinh r)$
$\quad\times(A_{3}\sinh r+(t_{2}+A_{4})\cosh r)\bigr]$
$\times\exp\bigl[(\tau_{2}\cosh r+A_{1}\sinh r)$
$\quad\times(A_{4}\sinh r+(\tau_{1}+A_{3})\cosh r)\bigr]\Bigr\}$
$=D_{1}(m_{1},n_{1},m_{2},n_{2}),$ (B5)
where
$\Gamma_{1}\{\cdot\}$ $=\frac{1}{N_{m,n}}\frac{\partial^{2m+2n+m_{1}+n_{1}+m_{2}+n_{2}}}{\partial t_{1}^{m}\partial\tau_{1}^{n}\partial t_{2}^{m}\partial\tau_{2}^{n}\partial x_{1}^{m_{1}}\partial y_{1}^{n_{1}}\partial x_{2}^{m_{2}}\partial y_{2}^{n_{2}}}\{\cdot\}$
$\quad\Big|_{t_{1}=\tau_{1}=t_{2}=\tau_{2}=x_{1}=y_{1}=x_{2}=y_{2}=0},$ (B6)
and
$A_{1}$ $=\frac{\sqrt{2}}{2}\bigl(\sqrt{1-l}e^{-i\phi}x_{1}+iy_{1}\bigr),$
$A_{2}$ $=\frac{\sqrt{2}}{2}\bigl(y_{2}-i\sqrt{1-l}e^{i\phi}x_{2}\bigr),$
$A_{3}$ $=\frac{\sqrt{2}}{2}\bigl(y_{1}+i\sqrt{1-l}e^{-i\phi}x_{1}\bigr),$
$A_{4}$ $=\frac{\sqrt{2}}{2}\bigl(\sqrt{1-l}e^{i\phi}x_{2}-iy_{2}\bigr).$ (B7)
Appendix C Specific Derivations and Expressions for $\langle I_{D}\rangle$ and $\langle I_{D}^{2}\rangle$ under $k=2$ Condition
With respect to operators that are partially obtained from the transformation relation of BS2, such as $a^{\dagger}b$, it is necessary to utilize the transformation relation of the Kerr nonlinear phase shift (Eq. (8)) and employ the IWOP technique to facilitate the calculation of the average values $\langle I_{D}\rangle$ and $\langle I_{D}^{2}\rangle$. Hence, we construct the general formula
$D_{2}(x)$
$={}_{f}\langle 0\rvert\langle\text{in}\rvert B_{1}^{\dagger}B_{f}^{\dagger}e^{-i\phi(a^{\dagger}a)^{2}}(a^{\dagger}b)^{x}e^{i\phi(a^{\dagger}a)^{2}}B_{f}B_{1}\lvert\text{in}\rangle\lvert 0\rangle_{f}$
$=\langle\psi\rvert e^{-x^{2}i\phi}a^{\dagger x}e^{-2xi\phi a^{\dagger}a}b^{x}\lvert\psi\rangle$
$=\frac{\partial^{x}}{\partial s^{x}}\langle\psi\rvert e^{-x^{2}i\phi}\exp[a^{\dagger}bs]e^{-2xi\phi a^{\dagger}a}\lvert\psi\rangle\Big|_{s=0},$ (C1)
where $x=1,2$ and for the convenience of using IWOP technology, $B_{f}B_{1}\lvert\text{in}\rangle\lvert 0\rangle_{f}$ is converted into a quantum state that represents a coherent state, denoted as $\lvert\psi\rangle$. Using the following normal ordering form
$e^{-2xi\phi a^{\dagger}a}=\colon\!\exp\bigl[(e^{-2xi\phi}-1)a^{\dagger}a\bigr]\!\colon$, (C2)
and integral formula
$\int\frac{d^{2}z}{\pi}e^{\zeta\lvert z\rvert^{2}+\xi z+\eta z^{\ast}+fz^{2}+gz^{\ast 2}}=\frac{e^{\frac{-\zeta\xi\eta+\xi^{2}g+\eta^{2}f}{\zeta^{2}-4fg}}}{\sqrt{\zeta^{2}-4fg}},$ (C3)
we can compute $D_{2}(x)$ as
$D_{2}(x)=\frac{\exp\bigl[-x^{2}i\phi-\alpha^{2}+M_{0}+\frac{-M_{1}M_{2}M_{3}+M_{4}M_{2}^{2}+M_{5}M_{3}^{2}}{M_{1}^{2}-4M_{4}M_{5}}\bigr]}{\cosh^{2}r\,N_{m,n}\sqrt{M_{1}^{2}-4M_{4}M_{5}}},$ (C4)
where
$M_{0}$ $=V_{5}\bigl(-i\tanh r(V_{1}+V_{2})(V_{1}-V_{2})V_{5}$
$\quad-i(V_{1}-V_{2})\tau_{1}+(V_{1}+V_{2})V_{4}\bigr),$
$M_{1}$ $=\tanh^{2}r\frac{4V_{1}^{2}+1+(1-l)s^{2}}{2}-1,$
$M_{2}$ $=\tau_{2}+\tanh r\bigl((V_{1}+V_{2})V_{4}-i(V_{1}-V_{2})\tau_{1}\bigr)$
$\quad-i\tanh^{2}r(V_{1}+V_{2})(V_{1}-V_{2})V_{5},$
$M_{3}$ $=i(V_{1}-V_{3})V_{4}+(V_{1}+V_{3})\tau_{1}$
$\quad+\tanh rV_{5}\frac{4V_{1}^{2}+1+(1-l)s^{2}}{2},$
$M_{4}$ $=i\tanh r(V_{1}+V_{3})(V_{1}-V_{3}),$
$M_{5}$ $=-i\tanh^{3}r(V_{1}+V_{2})(V_{1}-V_{2}),$ (C5)
where
$V_{1}$ $=\frac{(1-l)e^{-2xi\phi}+l}{2},$
$V_{2}$ $=\frac{1+i\sqrt{1-l}s}{2},$
$V_{3}$ $=\frac{1-i\sqrt{1-l}s}{2},$
$V_{4}$ $=t_{1}+\frac{\alpha}{\cosh r},$
$V_{5}$ $=t_{2}+\frac{\alpha}{\cosh r}.$ (C6)
By further utilizing Eqs. (6), (B1) and (C1), the expressions for $\langle I_{D}\rangle$ and $\langle I_{D}^{2}\rangle$ can be obtained for $k=2$ as
$\langle I_{D}\rangle=i\bigl[D_{2}(1)-D_{2}(1)^{\ast}\bigr],$ (C7)
and
$\langle I_{D}^{2}\rangle$ $=D_{1}(0,1,0,1)+D_{1}(1,0,1,0)$
$\quad+2D_{1}(1,1,1,1)-D_{2}(2)-D_{2}(2)^{\ast}.$ (C8)
Appendix D $C_{Q}$ for $k=2$ Case
In this appendix, we derive $C_{Q}$ for the Kerr nonlinear phase shift with $k=2$ and its specific expression to obtain the QFI $F_{L_{2}}$ under photon losses condition. To compute $C_{Q}$ using Eqs. (21), (22), (25) and (26) according to the method in Refs. 40 and 49 , we first obtain the normal ordering form of $(1-l)^{n_{a}}n_{a}^{q}$ by utilizing the operator identity from Eq. (16) as follows
$(1-l)^{n_{a}}n_{a}^{q}$ $=\eta^{n_{a}}n_{a}^{q}$
$=\frac{\partial^{q}}{\partial x^{q}}\exp\bigl[n_{a}\ln\eta\bigr]\exp\bigl[n_{a}x\bigr]\Big|_{x=0}$
$=\colon\!\frac{\partial^{q}}{\partial x^{q}}e^{(\eta e^{x}-1)a^{\dagger}a}\Big|_{x=0}\!\colon$, (D1)
where for simplicity, we set $\eta=1-l$. Based on this equation and using IWOP technology, the following summation expression $S_{q,p}$ can be further calculated for the operators associated with the Hermitian operators $H_{1,2}$ from Eqs. (21) and (22):
$S_{q,p}$ $=\sum_{j=0}^{\infty}\frac{(1-\eta)^{j}}{j!}j^{p}a^{\dagger j}\eta^{n_{a}}n_{a}^{q}a^{j}$
$=\sum_{j=0}^{\infty}\frac{(1-\eta)^{j}}{j!}j^{p}\colon\!(a^{\dagger}a)^{j}\frac{\partial^{q}}{\partial x^{q}}e^{(\eta e^{x}-1)a^{\dagger}a}\Big|_{x=0}\!\colon$
$=\colon\!\sum_{j=0}^{\infty}\frac{[(1-\eta)a^{\dagger}a]^{j}}{j!}\frac{\partial^{q+p}}{\partial x^{q}\partial y^{p}}e^{(\eta e^{x}-1)a^{\dagger}a+yj}\Big|_{x=y=0}\!\colon$
$=\colon\!\frac{\partial^{q+p}}{\partial x^{q}\partial y^{p}}e^{[\eta e^{x}+(1-\eta)e^{y}-1]a^{\dagger}a}\Big|_{x=y=0}\!\colon$
$=\frac{\partial^{q+p}}{\partial x^{q}\partial y^{p}}\bigl[\eta e^{x}+(1-\eta)e^{y}\bigr]^{n_{a}}\Big|_{x=y=0}.$ (D2)
The final step in the above equation employs the operator identity from Eq. (16) for conversion.
Similar to the calculation method for $C_{Q}$ in Eq. (20), substituting the Kraus operator with $k=2$ (Eq. (25)) into Eqs. (21) and (22), and combining with Eq. (D2) and the following transformation relationships,
$e^{\lambda a^{\dagger}a}a^{j}e^{-\lambda a^{\dagger}a}$ $=e^{-\lambda j}a^{j},$
$e^{\lambda(a^{\dagger}a)^{2}}a^{j}e^{-\lambda(a^{\dagger}a)^{2}}$ $=e^{\lambda j^{2}}a^{j}e^{-2\lambda ja^{\dagger}a},$ (D3)
we can further compute to obtain
$C_{Q}$ $=4\bigl[K_{1}^{2}\langle\Delta^{2}n_{a}^{2}\rangle_{S}-K_{2}\langle n_{a}^{3}\rangle_{S}+K_{3}\langle n_{a}^{2}\rangle_{S}$
$\quad-K_{4}\langle n_{a}\rangle_{S}-K_{5}\langle n_{a}^{2}\rangle_{S}\langle n_{a}\rangle_{S}-K_{6}\langle n_{a}\rangle_{S}^{2}\bigr],$ (D4)
where
$K_{1}$ $=\omega_{1}\eta^{2}-2\omega_{2}\eta-\mu_{2},$
$K_{2}$ $=2\eta\bigl[3\omega_{1}^{2}\eta^{3}-3\omega_{3}^{2}\eta^{2}-\omega_{4}\eta+\omega_{5}\bigr],$
$K_{3}$ $=\eta\bigl[11\omega_{1}^{2}\eta^{3}-2\omega_{6}^{2}\eta^{2}+\omega_{7}\eta-4\omega_{1}\omega_{2}\bigr],$
$K_{4}$ $=\eta\omega_{1}^{2}(6\eta^{3}-12\eta^{2}+7\eta-1),$
$K_{5}$ $=2(1-\eta)\eta\omega_{1}K_{1},$
$K_{6}$ $=(1-\eta)^{2}\eta^{2}\omega_{1}^{2},$ (D5)
and
$\omega_{1}$ $=1+2\mu_{1}-\mu_{2},$
$\omega_{2}$ $=\mu_{1}-\mu_{2},$
$\omega_{3}$ $=1+2(3\mu_{1}-2\mu_{2})+(2\mu_{1}-\mu_{2})(4\mu_{1}-3\mu_{2}),$
$\omega_{4}$ $=7\mu_{2}-6\mu_{1}+24\mu_{1}\mu_{2}-14\mu_{1}^{2}-9\mu_{2}^{2},$
$\omega_{5}$ $=\mu_{2}\omega_{1}-2\omega_{2}^{2},$
$\omega_{6}$ $=9+40\mu_{1}-22\mu_{2}+44\mu_{1}^{2}-48\mu_{1}\mu_{2}+13\mu_{2}^{2},$
$\omega_{7}$ $=7+40\mu_{1}-26\mu_{2}+52\mu_{1}^{2}-64\mu_{1}\mu_{2}+19\mu_{2}^{2},$ (D6)
where parameters $\mu_{1}$ and $\mu_{2}$ are optimizable to describe the photon losses occurring before and after the phase shifter. In particular, $\mu_{1}=\mu_{2}=0$ or 1-1 represents photon losses occurring before or after the phase shifter, respectively, as depicted in Fig. 7. Specifically, for the case $\mu_{1}=\mu_{2}=-1$, $F_{L_{2}}\leqslant C_{Q}=4\langle\Delta^{2}n_{a}^{2}\rangle$, representing the ideal case of QFI. Using Eqs. (D4)–(D6), optimizing through $\partial C_{Q}/\partial\mu_{1}=\partial C_{Q}/\partial\mu_{2}=0$ to find the minimum value of $C_{Q}$ gives
$\mu_{1\text{opt}}$ $=\frac{G_{2}G_{5}-G_{3}G_{4}}{G_{1}G_{4}-2\eta G_{2}^{2}},$ (D7)
$\mu_{2\text{opt}}$ $=\frac{G_{1}G_{5}-2\eta G_{2}G_{3}}{G_{1}G_{4}-2\eta G_{2}^{2}},$ (D8)
where
$G_{1}$ $=2\bigl[-(1-\eta)\eta\bigl(\langle\Delta^{2}n_{a}^{2}\rangle+2\langle n_{a}^{2}\rangle\langle n_{a}\rangle-\langle n_{a}\rangle^{2}\bigr)$
$\quad-(6\eta^{2}-6\eta+1)\bigl(\langle n_{a}^{3}\rangle+\langle n_{a}\rangle\bigr)$
$\quad+(11\eta^{2}-11\eta+2)\langle n_{a}^{2}\rangle\bigr],$
$G_{2}$ $=(1-\eta)^{2}\langle\Delta^{2}n_{a}^{2}\rangle+3(1-\eta)(2\eta-1)\langle n_{a}^{3}\rangle$
$\quad+(11\eta^{2}-13\eta+3)\langle n_{a}^{2}\rangle-(6\eta^{2}-6\eta+1)\langle n_{a}\rangle$
$\quad-(1-\eta)(2\eta-1)\langle n_{a}^{2}\rangle\langle n_{a}\rangle+\eta(1-\eta)\langle n_{a}\rangle^{2},$ (D9)
and
$G_{3}$ $=\eta^{2}\langle\Delta^{2}n_{a}^{2}\rangle-3\eta(2\eta-1)\langle n_{a}^{3}\rangle$
$\quad+(11\eta^{2}-9\eta+1)\langle n_{a}^{2}\rangle-(6\eta^{2}-6\eta+1)\langle n_{a}\rangle$
$\quad+\eta(2\eta-1)\langle n_{a}^{2}\rangle\langle n_{a}\rangle+\eta(1-\eta)\langle n_{a}\rangle^{2},$
$G_{4}$ $=-(1-\eta)^{3}\langle\Delta^{2}n_{a}^{2}\rangle-6\eta(1-\eta)^{2}\langle n_{a}^{3}\rangle$
$\quad-\eta(1-\eta)(11\eta-4)\langle n_{a}^{2}\rangle-\eta(6\eta^{2}-6\eta+1)\langle n_{a}\rangle$
$\quad+2\eta(1-\eta)^{2}\langle n_{a}^{2}\rangle\langle n_{a}\rangle+\eta^{2}(1-\eta)\langle n_{a}\rangle^{2},$
$G_{5}$ $=\eta\bigl[-\eta(1-\eta)\bigl(\langle\Delta^{2}n_{a}^{2}\rangle-\langle n_{a}\rangle^{2}\bigr)$
$\quad-(6\eta^{2}-6\eta+1)\bigl(\langle n_{a}^{3}\rangle+\langle n_{a}\rangle\bigr)$
$\quad+(11\eta^{2}-11\eta+2)\langle n_{a}^{2}\rangle$
$\quad+(2\eta^{2}-2\eta+1)\langle n_{a}^{2}\rangle\langle n_{a}\rangle\bigr],$ (D10)
and substituting these into Eq. (D4), along with using Eq. (16) for $w=1,2,3,4$, results in $F_{L_{2}}=C_{Q\min}$.